OpenStudy (mathmath333):

find \(\large\tt \color{black}{7^{123} (mod~~ 10)}\) using eulers theorm

OpenStudy (accessdenied):

Have you tried anything with this so far? I am not completely familiar with this but I think I found a workable solution...

OpenStudy (mathmath333):

i can find in general way ,i dont know how to use euler theorm

OpenStudy (accessdenied):

The statement of the theorem is that: If n and a are coprime, then \( a^{\phi(n)} \equiv 1 \ (\text{mod} \ n) \) What I am thinking is that, in our problem we are working with mod 10, so n = 10. And a = 7 clearly. So 7 and 10 are coprime, and the theorem applies. Now, \( \phi (10) = \phi (2*5) = (2 - 1)(5 - 1) = 4 \), and we have: \(7^{4} \equiv 1 \ (\text{mod} \ 10) \) Does that all make sense?

OpenStudy (accessdenied):

From that point, I figure we can rewrite \(7^{123} = 7^{4n + r} = 7^{4n} 7^r \), and cancel out the \(7^{4n} \) part as multiplication by 1. The remaining \(7^r\) should be much easier to take the mod of!

OpenStudy (mathmath333):

ok 7 and 10 are co prime ,how did u find that empty set (10), what if n=11

OpenStudy (mathmath333):

empty set=11-1=10 ?

OpenStudy (accessdenied):

Oh, \( \phi (n) \) is meant to be the Euler totient function. Basically it is a count of all positive numbers less than n that are relatively prime to n. You could count them directly: 1, 3, 7, 9 (so phi(10) = 4). That trick I used I saw earlier and is basically using two facts: \( \phi (nm) = \phi (n) \phi (m)\) and \(\phi (p) = p - 1\) for a prime p.

OpenStudy (accessdenied):

I have to go for now, sorry! There is a good example similar to this problem on the wiki page. "The theorem may be used to easily reduce large powers modulo n. ..." http://en.wikipedia.org/wiki/Euler's_theorem And more about the Euler totient function: http://en.wikipedia.org/wiki/Euler's_totient_function If you need more assistance, definitely bump the question! Good luck!

OpenStudy (mathmath333):

i got the formula of emptyset(n) for any n|N \(\large\tt \color{black}{\large\tt \color{black}{\emptyset(n)=n[(1-\dfrac{1}{x})(1-\dfrac{1}{y}......]}}\) where x ,y,,,,, are prime factors

OpenStudy (mathmath333):

thanks very much for your help

OpenStudy (accessdenied):

It isn't actually the empty set symbol, this is the greek letter phi: \( \phi \). It represents the Euler totient function or Euler phi function. Although in LaTeX it defaults to that display with \phi, but it can also show up as this: \( \varphi \) using \varphi. Maybe that's what mixed you up there! If I use that phi: \( \large a^{\varphi (n)} \equiv 1 \ (\text{mod} \ n) \)

OpenStudy (mathmath333):

ok thanks though

OpenStudy (accessdenied):

Glad to help! :)